starbucksgirl13:

Show your work and explain in a full sentence how 4 2/6 is equivalent to 3 8/6 .

starbucksgirl13:

plzzzzzzz help

Timmyspu:

ok so first what number can you divide to get the other fractions

Timmyspu:

Can you help me @Isry

KAPALOTJAY:

The given fractions that we are comparing are expressed as mixed numbers. It means that each is made up of whole numbers and fractions. We would convert each mixed number to improper fraction. By converting to improper fraction, we would multiply the whole number by the denominator and add the product to the numerator. The fraction would be the ratio of the result to the denominator. Considering 4 2/6, it becomes (4 × 6) + 2) = 26/6 Considering 3 8/6, it becomes (3 × 6) + 8)/6 = 26/6 Therefore, they are equivalent

starbucksgirl13:

ok thx

KAPALOTJAY:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @starbucksgirl13 ok thx \(\color{#0cbb34}{\text{End of Quote}}\) yw medal?

starbucksgirl13:

okkkkk